The two-year interest rate of 30 3⁄4% is equal to what nominal rate of interest convertible quarterly?
Answer #1

Hi

two year Effective interest rate EAR= 30 3/4 %

= 123/4 %

nominal quarterly rate = (1+EAR)^(1/n)-1

Here n = 8 since in 2 years, there are 8 quarters

nominal quarterly rate = (1+123/4%)^(1/8)-1

= 1.3075^(1/8)-1

= 0.034083

=3.41%

Thanks